Identify the range of the function:
y = 2/3x + 1/3 domain: 4, 6, 8 ,12

Answer by mananth(16946)


Identify the range of the function:
y = 2/3x + 1/3 domain: 4, 6, 8 ,12
x= 4
y = (2/3)*4+(1/3) = 9/3 = 3
x= 6
y = (2/3)*6+(1/3) = `13/3
x=8
x = (2/3)*8+(1/3) = `17/3
x = 12
(2/3)*12+(1/3) = `25/3
herefore, the range of the function for y = 2/3x + 1/3 for the given domain (4, 6, 8, 12) is
3, (13/3), (17/3), (25/3) respectively